What is a perimeter of a rectangle that measuers 7 feet by 4 feet?
Answer by ReadingBoosters(3246)

Perimeter is the distance around...
7+4+7+4 = 22
In other words, p = 2(l)+2(w)
therefore, p = 2(7) + 2(4) = 14 + 8 = 22
or, p = 2(l+w)
p = 2(7+4) = 2(11) = 22
.....................
Have more questions?
Email: HomeworkHelpers@readingboosters.com
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.